calculate urms of molecules of H2 at 1 atmp if density of H2 is 0.00009g/cc

Grade:11

1 Answers

Subhash Kumar Mehta
6 years ago
Density=0.00009g/cc=0.09kg/mcPressure=1 atm=1.013×10^5 N/m^3So Vrms=sq root(3P/p),where p=densityVrms=sq root(3×1.013×10^5/0.09) =1837.57If you like plz approve the answer